### v1

# Dinkel Yan Wang Wang Sun Niu Liu Li Zhang Luan MiLM Plus, Xiaomi Inc.China

Abstract
Contrastive Language Audio Pretraining (CLAP) is a widely-used method to bridge the gap between audio and text domains. Current CLAP methods enable sound and music retrieval in English, ignoring multilingual spoken content. To address this, we introduce general language audio pretraining (GLAP), which expands CLAP with multilingual and multi-domain abilities. GLAP demonstrates its versatility by achieving competitive performance on standard audio-text retrieval benchmarks like Clotho and AudioCaps, while significantly surpassing existing methods in speech retrieval and classification tasks. Additionally, GLAP achieves strong results on widely used sound-event zero-shot benchmarks, while simultaneously outperforming previous methods on speech content benchmarks. Further keyword spotting evaluations across 50 languages emphasize GLAP’s advanced multilingual capabilities. Finally, multilingual sound and music understanding is evaluated across four languages.
keywords:
contrastive language-audio pretraining, general pretraining, general audio encoders, large-language models
1 Introduction
In the field of computer vision, Contrastive Language-Image Pretraining (CLIP) [1] represents a significant breakthrough in extracting efficient representations that can be applied across various downstream tasks and domains. Similarly, Contrastive Language-Audio Pretraining (CLAP) [2, 3, 4] bridges text and audio domains, enabling zero-shot transfer learning i.e., testing the model on novel concepts that is has not seen during training. Notably, [5] trained on 4.6 Million pairs of audio and speech data, but has shown poor results for trivial speech classification tasks such as keyword spotting (see MSCLAP-2023 in Figure 1). While multilingual extensions [6] improved retrieval performance across eight languages, their approach still lacks basic speech understanding. CLAP embeddings primarily target sound and music, missing comprehensive speech representation (i.e., spoken language) - a critical aspect of audio processing. While there has been previous work focusing on speech-text embeddings using contrastive learning [3], a general approach that can be used between sound, music and speech domains is still missing. This work proposes general language audio pretraining (GLAP), an extension of previous CLAP works, aimed at aligning speech content with text, without compromising in sound and music performance. Our experiments demonstrate that GLAP achieves competitive performance in music and sound retrieval tasks while significantly improving speech understanding capabilities. GLAP also effectively generalizes its speech and sound understanding capabilities beyond English.
Figure 1: GLAP’s retrieval and zero-shot performance. A@T and T@A represent retrieval tasks of Audio-to-Text and Text-to-Audio, respectively, others are zero-shot (number of labels in brackets). Missing baselines were evaluated by the authors.
2 General language audio pretraining
To enable speech understanding in CLAP models, which are trained on sound and music data, one simple solution is to add speech data to the training dataset. However, as our analysis in Section 4.1 shows, this approach leads to compromised performance due to the absence of a unified audio encoder. The model either performs well on sound/music or on speech, but struggles to excel at both simultaneously. Thus, GLAP has two primary goals:
  1. 1.
    Deliver a unified encoder framework that maintains high performance for sound, music and speech retrieval tasks to enable alignment across these audio modalities.
  2. 2.
    Enable multilingual search capabilities for sound, music and speech content.
GLAP is trained with pairs of audio-text samples $(a,t)$ in contrastive fashion. Features from these audio-text pairs are extracted through a pre-trained multi-lingual text encoder ${\text{E}_{T}}$ and a pre-trained general audio encoder ${\text{E}_{A}}$:
$e_{a}$ $=\text{MLP}_{A}(\text{E}_{A}(a)),$ $e_{t}=\text{MLP}_{T}(\text{E}_{T}(t)),$
A trainable multi-layer perceptron (MLP) is added to align the dimensions. Finally, the pair $(e_{a},e_{t})$ is scored using cosine distance $s=\frac{e_{a}\cdot e_{t}^{T}}{||e_{a}||\cdot||e_{t}||}$. Unlike previous works, we use the sigmoid loss [7] as our main training objective $\mathcal{L}$, computed as:
$\mathcal{L}$ $=-\frac{1}{B}\sum_{i}^{B}\sum_{j}^{B}\log\sigma\left(s^{\prime}(i,j)\cdot\psi[i,j]\right),$ (1)
$s^{\prime}(i,j)$ $=\frac{s(i,j)+\beta}{\tau},$ (2)
where $\sigma$ is the sigmoid function, $B$ is the batchsize, $\beta,\tau$ are learnable parameters, $\cdot$” is the element-wise product, and
$\psi[i,j]=\begin{cases}1&\text{if }i=j,\\ -1&\text{otherwise}.\end{cases}$
The primary reason for choosing sigmoid loss over standard cross-entropy is its superior performance with large batch sizes and datasets, as we observed performance boosts of 1% to 5% across all retrieval tasks. An overview of the proposed framework can be seen in Figure 2.
Figure 2: The GLAP framework. GLAP is trained via contrastive learning where positive pairs ($\psi[i,j]=1$) are shown in blue and negative pairs ($\psi[i,j]=-1$) in white, with the added benefits of enabling multilingual speech-content retrieval, ontop of the standard sound/music capabilities.
3 Experimental Setup
The audio data is preprocessed by resampling all datasets to a single channel at 16 kHz. The trainable loss parameters $\tau,\beta$ (Equation 2) are initialized as $0.07$ and $-10$ respectively. For text embeddings, we use the text encoder from Sonar [8] as the default model, following [6]. We use a batch size of 128 per GPU across eight A800 GPUs, resulting in an effective batch size of $B=1024$. Embeddings are gathered across all GPUs before calculating the loss. Each epoch is defined as processing 10,000 batches, with training running for a maximum of 20 epochs. Model training employs an 8-bit Adam optimizer with a cosine decay scheduler. The learning rate starts at 0, warms up to ${10^{-4}}$ over the first two epochs, and decays to ${10^{-5}}$ over the remaining training period. Training takes approximately 1.5 days, with the best models typically achieved within the first 15 epochs. The source code and checkpoints are publicly available111github.com/xiaomi-research/dasheng-glap.
3.1 Datasets
Training
In this work, we integrate a wide range of existing audio-text datasets, as outlined in Table 1. Our primary speech training dataset is YODAS [9], a 400k-hour YouTube corpus labeled mainly via automated speech-to-text pipelines. Due to its noisy labeling, we supplement it with cleaner English (GigaSpeech, LibriSpeech) and Chinese (AISHELL-1) datasets. While YODAS covers 145 languages, relying solely on multilingual speech data limits generalization in sound/music retrieval. To address this, we followed [6] and leveraged Sonar [8] to translate the original English captions of all sound and music datasets into other seven widely spoken languages: German, Chinese, Catalan, Spanish, Japanese, French, and Dutch. As shown in Table 1, the training data is skewed towards speech, leading to poor performance on non-speech tasks. To balance this, we categorize the data into four groups: sound + music, English speech (GigaSpeech + LibriSpeech + YODAS English), Chinese speech (AISHELL-1 + YODAS Chinese), and other languages in YODAS. During training, we sample equally from each group, ensuring a balanced training process.
Domain Dataset hours #Pairs #Lang
Speech YODAS [9] 400 k 431 M 145
GigaSpeech [10] 10 k 8 M 1
LibriSpeech [11] 960 271 k 1
AISHELL-1 [12] 180 131 k 1
Sound Sound-VECapsA [13] 5200 1.6 M 1+7
Auto-ACD [14] 5200 1.8 M 1+7
AudiosetCaps [15] 5700 2.0 M 1+7
WavCaps [16] 7544 400 k 1+7
AudioCaps [17] 127 49 k 1+7
Clothov2 [18] 35 4884 1+7
Music MusicCaps [19] 7.3 2640 1+7
Songdescriber [20] 12 360 1+7
Table 1: Training datasets with duration (hours), audio-text pairs (# Pairs), and languages (# Lang). Music and sound data, labeled in English, are auto-translated into seven other languages via Sonar.
LibriSpeechOther MusicCaps AISHELL2-Test
Text-to-Audio Audio-to-Text Text-to-Audio Audio-to-Text Text-to-Audio Audio-to-Text
R@1 R@10 R@1 R@10 R@5 R@10 R@5 R@10 R@1 R@10 R@1 R@10
MSCLAP-2022 [2] 0.1 0.6 0.0 0.4 4.3 7.2 5.2 7.4 0.0 0.2 0 0.18
MSCLAP-2023 [5] 0.1 0.4 0.1 0.2 14.4 21.7 17.7 25.9 0.1 0.2 0.0 0.2
L-CLAP [3] 0.1 0.8 0.1 0.5 17.2 25.5 22.0 31.1 0 0.2 0.0 0.2
[…]
COLLAP-Roberta [21] - - - - 15.2 - 9.5 - - - - -
COLLAP-GPT [21] - - - - 17.4 - 10.3 - - - - -
BLAT [4] 0.0 0.8 0 0.4 3.2 5.1 3.9 5.8 0.0 0.2 0.0 0.2
M2D-Clap [22] 0.1 0.6 0.0 0.4 4.3 7.2 5.2 7.4 0.1 0.2 0.1 0.3
GLAP 93.8 96.8 91.8 94.4 30.3 41.2 15.0 44.4 98.5 99.7 99.1 99.7
Table 2: Retrieval results for music and speech datasets. indicates evaluation from a public checkpoint. Best in bold; higher is better.
Evaluation
We evaluate retrieval performance across seven test sets, including sound datasets such as Auto-ACD (ACD)[14], AudioCaps (AC)[17], Clothov2 (Clotho) [18], music datasets such as MusicCaps (MC) [19] and the speech datasets LibriSpeech (LS) [11], GigaSpeech [10] and AISHELL-2 (AIS2) [23]. For zero-shot classification, we primarily follow [5] and evaluate on ESC-50, FSD50K, UrbanSound8K (US8K), CREMA-D, GTZAN, NSynth instruments, Beijing-Opera, VocalSound, as well as Speech Commands V1/V2 (SCV1/2) [24] and Fluent Speech Commands (FSC) [25]. All test datasets with the exception of AIS2 are labeled in English.
3.2 Evaluation metrics
In audio-text retrieval tasks, performance is evaluated using recall at rank (R@k), where R@k is 1 if the target item appears in the top k retrieved items, otherwise 0. We also use mean average precision at rank 10 (mAP10) for a more comprehensive comparison. For zero-shot inference, accuracy is used for single-class classification, and mean average precision (mAP) is used for multi-label classification, as is standard practice [5].
3.3 Prompting
GLAP supports zero-shot inference, allowing the model to generate outputs directly from text prompts without prior training on specific tasks. In zero-shot scenarios, crafting an effective text prompt is crucial for achieving optimal performance. This is particularly important for GLAP, as it lacks a dedicated token to distinguish between a spoken word (e.g., “cat”) and a sound event (e.g., “the sound of a cat”). Prompts used in this work are depicted in Table 3.
Task Prompt
Speech {label}
Music The music in the style of {label}.
Sound The sound of {label} can be heard.
Table 3: Prompts for zero-shot evaluation.
4 Results
4.1 Audioencoder investigation
Given the extensive range of previously explored CLAP methods, one might question why earlier approaches failed to achieve strong performance in general audio-text pretraining. In our view, a key limitation lies in the reliance on sound-event audio encoders for CLAP, which we believe represents a significant bottleneck for performance. In this section, we examine the role of audio encoders in general contrastive audio-text learning. For this purpose, we compare the proposed training framework using five different audio-encoders, being Dasheng [26], CED-Base [27], Beats [28], Whisper-Base [29] and WavLM [30]. Each encoder is selected for its approximately 90M parameters and support for variable-length inputs, a crucial requirement for processing speech recognition datasets—a capability lacking in many CLAP audio encoders [2, 5, 3]. Each encoder is initialized from a publicly available checkpoint and trained on the same dataset described in Section 3.1, using the previously described settings (Section 3), with Sonar serving as the default text encoder.
As it can be seen in our results in Table 4, the choice of audio-encoder is vital to achieve a well-balanced performance across tasks. Sound-event encoders, such as CED and Beats, perform well in sound and music retrieval tasks but struggle with speech-related tasks. Conversely, Whisper and WavLM excel in speech-related retrieval but underperform in sound event and music datasets. Dasheng, on the other hand, proves to be the most versatile choice for general audio encoding, achieving competitive performance across sound, music, and speech domains. Based on these findings, all subsequent experiments utilize a single GLAP model with Dasheng as its audio encoder, without fine-tuning on target datasets.
Encoder Sound Music Speech
AC ACD MC LS-other AIS2
CED-Base 58.6 62.0 25.1 87.8 70.6
Beats 55.1 64.3 23.9 91.8 44.0
Whisper-Base 46.5 52.9 15.8 98.9 99.4
WavLM 36.1 47.5 14.8 99.9 96.3
Dasheng 55.8 60.1 20.3 94.8 99.0
Table 4: Text-to-Audio retrieval performance across five datasets, categorized by domain. LS-other refers to the test-other subset of LibriSpeech, while AIS2 corresponds to the AISHELL-2 test set. All experiments were conducted using the proposed training dataset and configuration. Values indicate mAP10, where higher scores represent better performance.
Method AudioCaps Clotho
Text-to-Audio Audio-to-Text Text-to-Audio Audio-to-Text
R@1 R@10 R@1 R@10 R@1 R@10 R@1 R@10
BLAT [4] 33.3 82.4 40.4 85.7 12.3 46.1 13.9 48.2
LClap-Large [3] 34.2 84.1 43.1 90.1 15.3 51.2 20.8 60.0
MSCLAP-2022 [2] 33.5 80.2 47.8 90.7 16.2 51.4 23.6 60.3
MSCLAP2023 [5] 35.6 - 42.5 - - 15.7 22.9 -
Wavcaps-CNN14 [16] 34.7 82.5 44.7 86.2 21.2 59.4 25.9 65.8
Wavcaps-HTSAT [16] 39.7 86.1 51.7 90.6 20.2 58.8 26.5 67.3
Auto-ACD [14] 39.5 85.4 53.7 91.7 15.3 52.1 17.7 52.6
T-CLAP [31] 39.7 86.9 49.8 91.9 17.3 53.6 21.8 57.4
MLCLAP [6] 40.7 87.8 50.1 92.8 18.8 59.0 21.1 62.5
Cacophony [32] 41.0 86.4 55.3 92.4 20.2 58.8 26.5 67.3
SoundVECaps [13] 41.2 85.3 53.3 93.0 - - - -
GLAP 41.7 86.1 54.4 91.1 19.4 58.3 21.8 61.5
Table 5: Sound event retrieval results compared to baselines.
Sound Music Speech
Method ESC50 FSD50K US8K CD VS GTZAN NS BO SCV1 SCV2 FSC
BLAT [4] 80.6 31.3 77.3 17.6 53.9 10.0 9.03 31.4 3.9 2.2 0.4
MS-CLAP-2023 [5] 88.2 40.3 75.0 29.7 69.2 58.4 47.9 46.6 16.4 2.5 0.3
[…]
L-CLAP [3] 91.0 21.5 77.0 18.3 79.3 47.4 26.1 40.2 3.8 4.1 0.3
M2D-Clap [22] 75.5 40.8 72.4 17.7 42.3 75.2 23.4 47.0 3.0 2.1 0.4
GLAP 88.8 40.9 78.9 20.5 75.1 69.6 31.3 36.5 96.6 95.8 75.6
Table 6: Zero-shot evaluation performance. Results marked with were obtained from a public checkpoint, while underlined entries used the corresponding training dataset and are therefore not truly zero-shot. Entries with used the 10-class variant instead of the 30-class version employed in this work. Best results are bolded and higher is better.
4.2 Sound retrieval results
The performance of GLAP on the widely used AudioCaps and Clotho datasets for English sound-event retrieval is presented in Table 5. GLAP demonstrates strong results on both benchmarks, surpassing other methods in Text-to-Audio retrieval (R@1) on AudioCaps while maintaining competitive performance on Clotho.
4.3 Music and speech retrieval results
For speech retrieval, we select the test-other dataset from Librispeech as a representative in-domain English speech benchmark, while the AISHELL-2 test set serves as an unseen Chinese evaluation set. To assess music-related capabilities, we use the widely adopted MusicCaps dataset [19]. Since some baseline models operate at different sampling rates (16 kHz vs. 44/48 kHz), we resample the test datasets accordingly. As shown in Table 2, our approach significantly outperforms previous methods in both music and speech retrieval. GLAP demonstrates exceptional multilingual retrieval performance, achieving over 93% on the English LibriSpeech-other test set and 98% on the Chinese AISHELL-2 test set. While most compared CLAP models offer a competitive performance on the MC dataset.
4.4 Zero shot evaluation
In this section, the zero-shot capabilities our approach are assessed using the prompts provided in Table 3. True zero-shot evaluation is difficult to determine, as some studies are trained on the respective datasets [3, 5] using its labels, meaning they are not strictly zero-shot. As shown in Table 6, our approach performs similarly to other baselines in sound-event and music classification. However, it excels in the keyword spotting task, significantly outperforming the baselines with accuracies of 95.8%, 96.6%, and 75.6% for the SCV1, V2, and FSC datasets, respectively. Notably, the FSC dataset requires understanding entire sentences, not just individual words, highlighting our model’s ability to align well with spoken content, as well as detecting single-word keywords (SCV1/2).
4.5 Multilingual capabilities
GLAP’s multilingual spoken content capabilities are assessed through a zero-shot evaluation on the Multilingual Spoken Words (MSW) Corpus [33].
Figure 3: Multilingual zero-shot keyword spotting (KWS) performance across 50 languages. Only the test set for each language is used, and the accuracies are reported. The number of keywords (num) for each language is shown on the right.
Results seen in Figure 3 show a strong zero-shot capability of GLAP on fifty languages. Oriya and Guarani achieve the best performance at 70% and 65.9% respectively. Notably, Chinese achieves 57.4% accuracy with 496 keywords, while Russian, containing 15,844 keywords, results in 39.6%. Additionally, we assess GLAP’s multilingual sound and music understanding by conducting zero-shot evaluations on the ESC-50, US8K and GTZAN datasets. For this, we use ChatGPT to translate the original (English) labels into the target language and adjust the prompts (Table 3) accordingly. As shown in Table 7, while performance drops compared to the English baseline, GLAP remains effective in multilingual sound-event classification. The model performs impressively in Russian (Ru) despite being trained only on Russian speech-text pairs from YODAS, not music/speech pairs (see Section 3.1). This suggests effective transfer of multilingual text-based knowledge to the audio domain.
Data Language
En De zh-CN Jp Ru
US8K 78.9 74.8 66.1 72.2 49.0
ESC-50 88.8 64.3 71.4 74.3 62.1
GTZAN 69.6 68.3 62.5 63.2 65.3
Table 7: GLAP’s zero-shot evaluation for multilingual sound and music. Original labels ( in gray) are translated into the target language using ChatGPT.
5 Conclusion
We introduce GLAP, a versatile language-audio pretraining framework that enables multilingual and multi-domain modeling of both audio and text. To the best of our knowledge, it is the first single system to integrate general audio and text embeddings into a unified contrastive framework. GLAP demonstrates competitive performance on well-established benchmarks like AudioCaps and Clotho, while surpassing previous methods in music and speech retrieval tasks. Zero-shot evaluations show strong results for English sound and music tasks, extending effectively to other languages. Inference on the Multilingual Spoken Words dataset highlights robust multilingual capabilities beyond English.