### v2

# GLAP: General contrastive audio-text pretraining across domains and languages

Abstract
Contrastive Language Audio Pretraining (CLAP) is a widely-used method to bridge the gap between audio and text domains. Current CLAP methods enable sound and music retrieval in English, ignoring multilingual spoken content. To address this, we introduce general language audio pretraining (GLAP), which expands CLAP with multilingual and multi-domain abilities. GLAP demonstrates its versatility by achieving competitive performance on standard audio-text retrieval benchmarks like Clotho and AudioCaps, while significantly surpassing existing methods in speech retrieval and classification tasks. Additionally, GLAP achieves strong results on widely used sound-event zero-shot benchmarks, while simultaneously outperforming previous methods on speech content benchmarks. Further keyword spotting evaluations across 50 languages emphasize GLAP’s advanced multilingual capabilities. Finally, multilingual sound and music understanding is evaluated across four languages.
Index Terms—  contrastive language-audio pretraining, general pretraining, general audio encoders, large-language models
1 Introduction
In the field of computer vision, Contrastive Language-Image Pretraining (CLIP) [21] represents a significant breakthrough in extracting efficient representations that can be applied across various downstream tasks and domains. Similarly, Contrastive Language-Audio Pretraining (CLAP) [25, 26] bridges text and audio, enabling zero-shot transfer learning. However, existing CLAP models, including those trained on large datasets like the one in [12] and multilingual extensions [27], perform poorly on basic speech tasks like keyword spotting. CLAP embeddings primarily target sound and music, missing comprehensive speech representation (i.e., spoken language) - a critical aspect of audio processing. This work proposes general language audio pretraining (GLAP) aimed at aligning speech content with text, without compromising in sound and music performance. GLAP further effectively generalizes its speech and sound understanding capabilities beyond English which be seen in Figure 1.
Fig. 1: GLAP’s retrieval and zero-shot performance. A@T and T@A represent retrieval tasks of Audio-to-Text and Text-to-Audio, respectively, others are zero-shot (#labels).
2 General language audio pretraining
To enable speech understanding in CLAP models, which are trained on sound and music data, one simple solution is to add speech data to the training dataset. However, as our analysis in Section 4.1 shows, this approach leads to compromised performance due to the absence of a unified audio encoder. Thus GLAP’s main goal is to provide a unified framework framework that maintains high performance for sound, music and speech retrieval tasks, enabling alignment across these audio modalities. GLAP is trained with pairs of audio-text samples $(a,t)$ in contrastive fashion. Features from these audio-text pairs are extracted through a pre-trained multi-lingual text encoder $\text{E}_{T}$} and a pre-trained general audio encoder $\text{E}_{A}$}:
$e_{a}$ $=\text{MLP}_{A}(\text{E}_{A}(a)),$ $e_{t}=\text{MLP}_{T}(\text{E}_{T}(t)),$
[…]
$\mathcal{L}$ $=-\frac{1}{B}\sum_{i}^{B}\sum_{j}^{B}\log\sigma\left(s^{\prime}(i,j)\cdot\psi[i,j]\right),$ (1)
$s^{\prime}(i,j)$ $=\frac{s(i,j)+\beta}{\tau},$ (2)
where $\sigma$ is the sigmoid function, $B$ is the batchsize, $\beta,\tau$ are learnable parameters, $\cdot$” is the element-wise product, and
$\psi[i,j]=\begin{cases}1&\text{if }i=j,\\ -1&\text{otherwise}.\end{cases}$
The primary reason for choosing sigmoid loss over standard cross-entropy is its superior performance with large batch sizes and datasets, as we observed performance boosts of 1% to 5% across all retrieval tasks. An overview of the proposed framework can be seen in Figure 2.
Fig. 2: GLAP enables multilingual speech-content retrieval, ontop of the standard sound/music capabilities.
3 Experimental Setup
All audio data is preprocessed by resampling all datasets to a single channel at 16 kHz. Following [27], we use the text encoder from Sonar [10] for text embeddings. We initialize the loss parameters as $\tau=0.07,\beta=-10$ (Equation 2). We train the model with an effective batch size of $B=1024$ (128 per GPU) for a maximum of 20 epochs (10,000 batches per epoch). Training uses an 8-bit Adam optimizer with a cosine decay scheduler, where the learning rate warms up from 0 to $10^{-4}$} over the first two epochs and then decays to $10^{-5}$}. The source code and checkpoints are publicly available111github.com/xiaomi-research/dasheng-glap.
Domain Dataset hours #Pairs #Lang
Speech YODAS [14] 400 k 431 M 145
GigaSpeech [4] 10 k 8 M 1
LibriSpeech [20] 960 271 k 1
AISHELL-1 [3] 180 131 k 1
Sound Sound-VECapsA [28] 5200 1.6 M 1+7
Auto-ACD [23] 5200 1.8 M 1+7
AudiosetCaps [2] 5700 2.0 M 1+7
WavCaps [18] 7544 400 k 1+7
AudioCaps [13] 127 49 k 1+7
Clothov2 [9] 35 4884 1+7
Music MusicCaps [1] 7.3 2640 1+7
Songdescriber [16] 12 360 1+7
Table 1: Training datasets with duration (hours), audio-text pairs (# Pairs), and languages (# Lang). Music and sound data labeled in English were auto-translated to 7 additional languages via Sonar.
LibriSpeechOther MusicCaps AISHELL2-Test
Text-to-Audio Audio-to-Text Text-to-Audio Audio-to-Text Text-to-Audio Audio-to-Text
R@1 R@10 R@1 R@10 R@5 R@10 R@5 R@10 R@1 R@10 R@1 R@10
MSCLAP-2022 [11] 0.1 0.6 0.0 0.4 4.3 7.2 5.2 7.4 0.0 0.2 0 0.18
MSCLAP-2023 [12] 0.1 0.4 0.1 0.2 14.4 21.7 17.7 25.9 0.1 0.2 0.0 0.2
L-CLAP [25] 0.1 0.8 0.1 0.5 17.2 25.5 22.0 31.1 0 0.2 0.0 0.2
[…]
COLLAP-Roberta [24] - - - - 15.2 - 9.5 - - - - -
COLLAP-GPT [24] - - - - 17.4 - 10.3 - - - - -
BLAT [26] 0.0 0.8 0 0.4 3.2 5.1 3.9 5.8 0.0 0.2 0.0 0.2
M2D-Clap [19] 0.1 0.6 0.0 0.4 4.3 7.2 5.2 7.4 0.1 0.2 0.1 0.3
GLAP 93.8 96.8 91.8 94.4 30.3 41.2 15.0 44.4 98.5 99.7 99.1 99.7
Table 2: Music/Speech retrieval results. indicates evaluation from a public checkpoint. Best in bold; higher is better.
3.1 Datasets
Training
We trained our model on a diverse range of audio-text datasets, as detailed in Table 1. Our main dataset is YODAS [14], a 400k-hour YouTube corpus primarily labeled with noisy automated speech-to-text pipelines. To improve data quality, we supplemented YODAS with cleaner English datasets (GigaSpeech, LibriSpeech) and a Chinese dataset (AISHELL-1). While YODAS covers 145 languages, relying solely on multilingual speech data limits generalization in sound/music retrieval. To address this, we followed [27] and leveraged Sonar [10] to translate the original English captions of all sound and music datasets into other seven widely spoken languages: German, Chinese, Catalan, Spanish, Japanese, French, and Dutch. To address the imbalance of speech data shown in Table 1, we categorized the data into four groups: sound + music, English speech, Chinese speech, and other languages and sample equally from each group.
Evaluation
We evaluate retrieval performance across sound datasets such as Auto-ACD (ACD) [23], AudioCaps (AC) [13], Clothov2 (Clotho) [9], music datasets such as MusicCaps (MC) [1] and the speech datasets LibriSpeech (LS) [20], GigaSpeech [4] and AISHELL-2 (AIS2) [3]. For zero-shot classification, we primarily follow [12] and evaluate on ESC-50, FSD50K, UrbanSound8K (US8K), CREMA-D, GTZAN, NSynth instruments, Beijing-Opera, VocalSound, as well as Speech Commands V1/V2 (SCV1/2) and Fluent Speech Commands (FSC) [15]. All test datasets with the exception of AIS2 are labeled in English.
3.2 Evaluation metrics and Zero-shot Prompting
Audio-text retrieval performance is measured using recall at rank (R@k), which checks if the correct item is in the top k results. For a more complete comparison, we also use mean average precision at rank 10 (mAP10). For zero-shot inference, we use accuracy for single-class classification and mean average precision (mAP) for multi-label classification, consistent with standard practice [12]. GLAP performs zero-shot inference, but without a dedicated token to differentiate between spoken words and sound events, the model relies on effective text prompts. Prompts used in this work are shown in Table 3.
Task Prompt
Speech {label}
Music The music in the style of {label}.
Sound The sound of {label} can be heard.
Table 3: Prompts for zero-shot evaluation.
4 Results
4.1 Audioencoder investigation
In our view, a key limitation lies in the reliance on sound-event audio encoders for CLAP. For this purpose, we compare the proposed training framework using five distinct audio-encoders, being Dasheng [8], CED-Base [7], Beats [6], Whisper-Base [22] and WavLM [5], each with $\approx$ 90M parameters and support for variable length inputs. Each encoder is initialized from a publicly available checkpoint and trained using the settings in Section 3. As the results in Table 4 indicate, sound-event encoders like CED and Beats excel at music and sound tasks, but perform poorly on speech. Conversely, Whisper and WavLM are strong on speech but weak on other audio types. Dasheng is the most versatile, performing well across all domains. Because of this, further experiments use Dasheng as the primary audio encoder.
Encoder Sound Music Speech
AC ACD MC LS-other AIS2
CED-Base 58.6 62.0 25.1 87.8 70.6
Beats 55.1 64.3 23.9 91.8 44.0
Whisper-Base 46.5 52.9 15.8 98.9 99.4
WavLM 36.1 47.5 14.8 99.9 96.3
Dasheng 55.8 60.1 20.3 94.8 99.0
Table 4: Text-to-Audio retrieval performance across five datasets in mAP10, where higher is better. LS-other represents test-other of LibriSpeech and AIS2 represents the AISHELL-2 test set.
Sound Music Speech
Method ESC50 FSD50K US8K CD VS GTZAN NS BO SCV1 SCV2 FSC
BLAT [26] 80.6 31.3 77.3 17.6 53.9 10.0 9.03 31.4 3.9 2.2 0.4
MS-CLAP-2023 [12] 88.2 40.3 75.0 29.7 69.2 58.4 47.9 46.6 16.4 2.5 0.3
[…]
L-CLAP [25] 91.0 21.5 77.0 18.3 79.3 47.4 26.1 40.2 3.8 4.1 0.3
M2D-Clap [19] 75.5 40.8 72.4 17.7 42.3 75.2 23.4 47.0 3.0 2.1 0.4
GLAP 88.8 40.9 78.9 20.5 75.1 69.6 31.3 36.5 96.6 95.8 75.6
Table 5: Zero-shot evaluation performance. Results marked with were obtained from a public checkpoint, while underlined entries used the corresponding training dataset and are therefore not truly zero-shot. Entries with used the 10-class variant instead of the 30-class version employed in this work. Best results are bolded and higher is better.
Method AudioCaps Clotho
Text-to-Audio Audio-to-Text Text-to-Audio Audio-to-Text
R@1 R@10 R@1 R@10 R@1 R@10 R@1 R@10
BLAT [26] 33.3 82.4 40.4 85.7 12.3 46.1 13.9 48.2
LClap-Large [25] 34.2 84.1 43.1 90.1 15.3 51.2 20.8 60.0
MSCLAP2023 [12] 35.6 - 42.5 - - 15.7 22.9 -
Wavcaps-CNN14 [18] 34.7 82.5 44.7 86.2 21.2 59.4 25.9 65.8
Wavcaps-HTSAT [18] 39.7 86.1 51.7 90.6 20.2 58.8 26.5 67.3
Auto-ACD [23] 39.5 85.4 53.7 91.7 15.3 52.1 17.7 52.6
MLCLAP [27] 40.7 87.8 50.1 92.8 18.8 59.0 21.1 62.5
SoundVECaps [28] 41.2 85.3 53.3 93.0 - - - -
GLAP 41.7 86.1 54.4 91.1 19.4 58.3 21.8 61.5
Table 6: Sound event retrieval results compared to baselines.
4.2 Sound retrieval results
The performance of GLAP on the widely used AudioCaps and Clotho datasets for English sound-event retrieval is presented in Table 6. GLAP demonstrates strong results on both benchmarks, surpassing other methods in Text-to-Audio retrieval (R@1) on AudioCaps while maintaining competitive performance on Clotho.
4.3 Music and speech retrieval results
We assessed GLAP’s retrieval capabilities using the Librispeech test-other dataset for English speech and the AISHELL-2 test set for unseen Chinese. For music, we used the MusicCaps dataset [1]. As shown in Table 2, GLAP significantly outperforms previous methods in both music and speech retrieval. It achieves over 93% on the English LibriSpeech test set and 98% on the Chinese AISHELL-2 test set, demonstrating exceptional multilingual performance.
4.4 Zero shot evaluation
Zero-shot results are shown in Table 5. Our model performs similarly to other baselines for sound and music classification. However, it significantly outperforms them in keyword spotting tasks, achieving 95.8%, 96.6%, and 75.6% accuracy on the SCV1, V2, and FSC datasets, respectively.
4.5 Multilingual capabilities
GLAP’s multilingual capabilities are evaluated using zero-shot tests on the Multilingual Spoken Words (MSW) Corpus [17].
Fig. 3: Multilingual zero-shot keyword spotting performance across 50 languages. The number of keywords (num) for each language is shown on the right.
Results in Figure 3 show strong performance across 50 languages. For example, Oriya and Guarani achieve the best accuracy at 70% and 65.9%, respectively. Notably, Chinese, with 496 keywords, reaches 57.4% accuracy, while Russian, with 15,844 keywords, achieves 39.6%. We also assessed GLAP’s multilingual sound and music understanding on the ESC-50, US8K, and GTZAN datasets. As shown in Table 7, while performance decreases compared to the English baseline, GLAP remains effective for multilingual sound classification.
Data Language
En De zh-CN Jp Ru
US8K 78.9 74.8 66.1 72.2 49.0
ESC-50 88.8 64.3 71.4 74.3 62.1
GTZAN 69.6 68.3 62.5 63.2 65.3
Table 7: GLAP’s zero-shot evaluation for multilingual sound and music. Original labels ( in gray) are translated into the target language using ChatGPT.
5 Conclusion
We introduced GLAP, a versatile language-audio pretraining framework that enables multilingual and multi-domain modeling of both audio and text. To the best of our knowledge, it is the first single system to integrate general audio and text embeddings into a unified contrastive framework. GLAP demonstrates competitive performance on well-established benchmarks like AudioCaps and Clotho, while surpassing previous methods in music and speech retrieval tasks. Zero-shot evaluations show strong results for English sound and music tasks, extending effectively to other languages. Inference on the Multilingual Spoken Words dataset highlights robust multilingual capabilities beyond English.